# Towards Efficient Training and Evaluation of Robust Models against $l_{0}$ Bounded Adversarial Perturbations

### v1

@@ A.6 More Ablation Study on sPGD
Table 11: Robust accuracy at different step sizes $\alpha$ for magnitude $p$. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ [19] trained on CIFAR-10 [59].
@@
Table 12: Robust accuracy at different step sizes $\beta$ for mask $m$. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ [19] trained on CIFAR-10 [59].
$\beta$ $2$ $4$ $8$ $16$ $24$ $32$
@@
A.7 Impact of Tolerance during Adversarial Training
We also study the impact of different tolerances for reinitialization during adversarial training. The results reported in Table A.7 indicate that higher tolerance during adversarial training significantly improves the model’s robustness against sparse attacks, and the performance reaches its best when the tolerance is set to $10$.
Table 13: Robust accuracy at different tolerance for reinitialization $t$ during attacking. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ [19] trained on CIFAR-10 [59].
@@ A.7 Impact of Tolerance during Adversarial Training
Appendix B Implementation Details
In experiments, we mainly focus on the cases of the sparsity of perturbations $k=10,15$ and $20$, where $k=||\sum_{i=1}^{c}\bm{\delta}^{(i)}||_{0}$ or $||{\bm{m}}||_{0}$, $\bm{\delta}^{(i)}\in{\mathbb{R}}^{h\times w}$ is the $i$-th channel of perturbation $\bm{\delta}\in{\mathbb{R}}^{h\times w\times c}$, and ${\bm{m}}\in{\mathbb{R}}^{h\times w\times 1}$ is the sparsity mask in the decomposition of $\bm{\delta}={\bm{p}}\odot{\bm{m}}$, ${\bm{p}}\in{\mathbb{R}}^{h\times w\times c}$ denotes the magnitude of perturbations.
sparse-PGD (sPGD): Cross-entropy loss is adopted as the loss function of both untargeted and targeted versions of our method. The step size for magnitude ${\bm{p}}$ is set $\alpha=1/4$; the step size for continuous mask $\widetilde{{\bm{m}}}$ is set $\beta=1/4\times\sqrt{h\times w}$, where $h$ and $w$ are the height and width of the input image ${\bm{x}}\in{\mathbb{R}}^{h\times w\times c}$, respectively. The small constant $\gamma$ to avoid numerical error is set to $1\times 10^{-10}$. The number of iterations is $10000$ for all datasets to ensure fair comparison among attacks in Table 1. The tolerance for reinitialization is set to $3$.
sparse-AutoAttack (sAA): It is a cascade ensemble of five different attacks, i.e., a) untargeted sPGD with unprojected gradient (sPGDunproj), b) untargeted sPGD with sparse gradient (sPGDproj), and c) untargeted Sparse-RS. The hyper-parameters of sPGD are the same as those listed in the last paragraph.
Adversarial Training: sPGD is adopted as the attack during the training phase, the number of iterations is $20$, and the backward function is randomly selected from the two different backward functions for each batch. For sTRADES, we only compute the TRADES loss when training, and generating adversarial examples is based on cross-entropy loss. We use PreactResNet18 [63] with softplus activation [66] for experiments on CIFAR-10 and CIFAR-100 [59], and ResNet34 [62] for experiments on ImageNet100 [60]. We train the model for $100$ epochs, the training batch size is $128$. The optimizer is SGD with a momentum factor of $0.9$ and weight decay factor of $5\times 10^{-4}$. The learning rate is initialized to $0.05$ and is divided by a factor of $10$ at the $30$th epoch and the $35$th epoch. The tolerance for reinitialization is set to $10$.
Appendix C Visualization of Some Adversarial Samples
We show some adversarial examples with different sparsity levels of perturbation in Figure 5, 6, 7. The attack is sPGD, and the model is Fast-EG-$l_{1}$ [19] trained on CIFAR-10. We can observe that most of the perturbed pixels are located in the foreground of images. It is consistent with the intuition that the foreground of an image contains most of the semantic information.
(a) automobile
(b) frog
(c) horse
(d) deer
(e) cat
(f) truck
(g) deer
(h) automobile
(i) bird
(j) ship
Figure 5: Clean images (first row) from the test set of CFIFAR-10 and their corresponding adversarial samples (second row) by sPGD. The attack is sPGD with sparsity level $k=10$. The model is Fast-EG-$l_{1}$ trained on CIFAR-10. The predictions given by the model are listed below the images.
(a) horse
(b) frog
(c) truck
(d) airplane
(e) bird
(f) cat
(g) bird
(h) horse
(i) automobile
(j) horse
Figure 6: Clean images (first row) from the test set of CFIFAR-10 and their corresponding adversarial samples (second row) by sPGD. The attack is sPGD with sparsity level $k=15$. The model is Fast-EG-$l_{1}$ trained on CIFAR-10. The predictions given by the model are listed below the images.
(a) dog
(b) bird
(c) cat
(d) horse
(e) ship
(f) cat
(g) automobile
(h) bird
(i) dog
(j) cat
Figure 7: Clean images (first row) from the test set of CFIFAR-10 and their corresponding adversarial samples (second row) by sPGD. The attack is sPGD with sparsity level $k=20$. The model is Fast-EG-$l_{1}$ trained on CIFAR-10. The predictions given by the model are listed below the images.

### v3

@@ Appendix C Ablation Studies
Table 10: Ablation study of each component in sPGDproj in terms of robust accuracy. The model is Fast-EG-$l_{1}$ trained on CIFAR-10.
Ablations Robust Acc.
Baseline (PGD0 w/o restart) 49.4
    + Decomposition: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ 58.0 (+8.6)
    + Continuous mask $\widetilde{{\bm{m}}}$ 33.9 (-15.5)
    + Random reinitialization 18.1 (-31.3)
sPGD: We first validate the effectiveness of each component in sPGD. The result is reported in Table 10. We observe that naively decomposing the perturbation $\bm{\delta}$ by $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ and updating them separately can deteriorate the performance. By contrast, the performance significantly improves when we update the mask ${\bm{m}}$ by its continuous alternative $\widetilde{{\bm{m}}}$ and $l_{0}$ ball projection. This indicates that introducing $\widetilde{{\bm{m}}}$ greatly mitigates the challenges in optimizing discrete variables. Moreover, the results in Table 10 indicate the performance can be further improved by the random reinitialization mechanism, which encourages exploration and avoids trapping in a local optimum.
In addition, we compare the performance when we use different step sizes for the magnitude tensor ${\bm{p}}$ and the sparsity mask ${\bm{m}}$. As shown in Table D and D, the robust accuracy does not vary significantly with different step sizes. It indicates the satisfying robustness of our method to different hyperparameter choices. In practice, We set $\alpha$ and $\beta$ to $0.25$ and $0.25\times\sqrt{hw}$, respectively. Note that $h$ and $w$ denote the height and width of the image, respectively, which are both $32$ in CIFAR-10.
Ultimately, we compare the performance of sPGD with different tolerance iterations for reinitialization. As shown in Table D, the performance of our method remains virtually unchanged, which showcases that our approach is robust to different choices of tolerance for reinitialization.
Adversarial training: We conduct preliminary exploration on adversarial training against sparse perturbations, since sPGD can be incorporated into any adversarial training variant. Table 1278 and 9 study sAT and sTRADES, while sTRADES outperforms sAT in all cases. In addition, the training of sAT is relatively unstable in practice, so we focus on sTRADES for ablation studies in this section. We leave the design of sPGD-adapted adversarial training variants to further improve model robustness as a future work.
Table D demonstrates the performance when we use different backward functions. The policies include always using the sparse gradient (Proj.), always using the unprojected gradient (Unproj.), alternatively using both backward functions every $5$ epochs (Alter.) and randomly selecting backward functions (Rand.). The results indicate that randomly selecting backward functions has the best performance. In addition, Table D demonstrates the robust accuracy of models trained by sTRADES with different multi-$\epsilon$ strategies. The results indicate that multi-$\epsilon$ strategy helps boost the performance. The best robust accuracy is obtained when the adversarial budget for training is $6$ times larger than that for test. Furthermore, we also study the impact of different tolerances during adversarial training in Table D. The results show that higher tolerance during adversarial training benefits the robustness of the model, and the performance reaches its best when the tolerance is set to $10$.
Structured Sparse Perturbations: Since the optimal hyperparameters for structured sparse perturbations differ from those for unstructured $l_{0}$ bounded perturbations, we conduct further ablation studies in the structured sparse setting.
The results in Table D, D and D suggest that the step size $\alpha$, $\beta$ and the tolerance $t$ for generating structured sparse perturbations should be smaller than those for unstructured cases. This can be attributed to the requirement for more exploitable attacks when generating structured sparse perturbations, similar to conventional $l_{\infty}$ and $l_{2}$ attacks [39].
The multi-$\epsilon$ strategy has been proven effective in $l_{0}$ adversarial training. However, in the context of structured sparse attack, particularly in the patch setting, we can increase either the sparsity level $\epsilon$ or the patch size to increase the corresponding $l_{0}$ adversarial budget. Therefore, we compare two different strategies here, i.e., multi-$\epsilon$ strategy and multi-size strategy. Specifically, the multi-$\epsilon$ strategy increases the sparsity level $\epsilon$ while maintaining the patch size unchanged during training, but the multi-size strategy only increases the patch size to make the $l_{0}$ sparsity level of generated perturbations similar to that in the multi-$\epsilon$ strategy. The results reported in Table D indicate that the multi-$\epsilon$ strategy provides better performance.
Appendix D More Visualization
More adversarial examples are shown in Figure 6. Additionally, we showcase some $l_{0}$ bounded adversarial examples generated on our sAT model in Figure 7. Compared to Figure 5 (a)-(b) and Figure 6 (a)-(b), the perturbations generated on the adversarially model are mostly located in the foreground of images. It is consistent with the intuition that the foreground of an image contains most of the semantic information [41].
Table 11: Robust accuracy at different step sizes $\alpha$ for magnitude $p$. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ trained on CIFAR-10.
@@ Appendix D More Visualization
Table 12: Robust accuracy at different step sizes $\beta$ for mask $m$. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ trained on CIFAR-10. Note that $h$ and $w$ are both $32$ in CIFAR-10.
$\beta/\sqrt{hw}$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ $\frac{3}{4}$ $1$
@@
Table 13: Robust accuracy at different tolerance for reinitialization $t$ during attacking. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ trained on CIFAR-10.
@@
Policy Proj. Unproj. Alter. Rand.
Acc. 41.5 39.4 51.5 61.7
Table 15: Ablation study on multi-$\epsilon$ strategy during adversarial training. The model is PRN18 trained by sTRADES with random policy. The robust accuracy is obtained through sAA.
$\epsilon$ 1×1$\times$1 × 2×2$\times$2 × 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 × 10×10$\times$10 ×
Acc. 34.0 39.7 54.5 61.7 60.2 55.5
@@
Table 17: Robust accuracy at different step sizes $\alpha$ for magnitude $p$. The evaluated attack is sPGDproj generating $5\times 5$ patch ($\epsilon=1$). The model is sTRADES-$l_{0}$trained on CIFAR-10.
$\alpha$ $0.01$ $0.0125$ $0.05$ $0.1$ $0.15$ $0.25$
Acc. 28.1 26.7 33.8 41.5 47.9 58.1
Table 18: Robust accuracy at different step sizes $\beta$ for mask $m$. The evaluated attack is sPGDproj generating $5\times 5$ patch ($\epsilon=1$). The model is sTRADES-$l_{0}$ trained on CIFAR-10. Note that $h$ and $w$ are both $32$ in CIFAR-10.
$\beta/\sqrt{hw}$ $0.01$ $0.0125$ $0.05$ $0.1$ $0.15$ $0.25$
Acc. 27.4 26.7 34.1 37.5 38.3 35.1
Table 19: Robust accuracy at different tolerance for reinitialization $t$ during attacking. The evaluated attack is sPGDproj generating $5\times 5$ patch ($\epsilon=1$). The model is sTRADES-$l_{0}$ trained on CIFAR-10.
$t$ 3 10 50 100 150
Acc. 36.1 28.7 26.7 28.1 33.7
Table 20: Ablation study on multi-$\epsilon$ / multi-size strategy during adversarial training against $3\times 3$ patch ($\epsilon=2$). The model is PRN18 trained by sTRADES with random policy.
Attack RS LOAP sPGDproj sPGDunproj sAA
Multi-$\epsilon$ 77.5 81.4 72.8 81.7 72.3
Multi-size 75.5 79.2 68.4 79.8 67.9
(a) $l_{0}$, $\epsilon=100$
(b) $l_{0}$, $\epsilon=100$
(c) column, $\epsilon=1$
(d) column, $\epsilon=1$
(e) patch, $14\times 14$, $\epsilon=1$
(f) patch, $14\times 14$, $\epsilon=1$
(g) patch, $10\times 10$, $\epsilon=2$
(h) patch, $10\times 10$, $\epsilon=2$
(i) heart, $20\times 20$, $\epsilon=1$
(j) heart, $20\times 20$, $\epsilon=1$
(k) circle, $60\times 60$, $\epsilon_{\infty}=64/255$
(l) circle, $60\times 60$, $\epsilon_{\infty}=64/255$
[…]
(a) $l_{0}$, $\epsilon=200$
(b) $l_{0}$, $\epsilon=200$
(c) $l_{0}$, $\epsilon=200$
(d) $l_{0}$, $\epsilon=200$
[…]